given the points below, find xy. round to the nearest hundredth. x(-9,2) and y(5,-4)


1
Expert's answer
2021-05-12T06:59:04-0400

"xy=\sqrt{(5-(-9))^2+(-4-2)^2}"


"xy=\sqrt{(5+9)^2+(-6)^2}"


"xy=\sqrt{(14)^2+(-6)^2}"


"xy=\sqrt{196+36}"


"xy=\sqrt{232}=15.2315462117"


"Hence,"

"xy=15.23"